### v1

# Dimensional reduction and adaptation-development-evolution relation in evolved biological systems

Abstract
Life systems are complex and hierarchical, with diverse components at different scales, yet they sustain themselves, grow, and evolve over time. How can a theory of such complex biological states be developed? Here we note that for a hierarchical biological system to be robust, it must achieve consistency between micro-scale (e.g. molecular) and macro-scale (e.g. cellular) phenomena. This allows for a universal theory of adaptive change in cells based on biological robustness and consistency between cellular growth and molecular replication. Here, we show how adaptive changes in high-dimensional phenotypes (biological states) are constrained to low-dimensional space, leading to the derivation of a macroscopic law for cellular states. The theory is then extended to evolution, leading to proportionality between evolutionary and environmental responses, as well as proportionality between phenotypic variances due to noise and due to genetic changes. The universality of the results across several models and experiments is demonstrated. Then, by further extending the theory of evolutionary dimensional reduction to multicellular systems, the relationship between multicellular development and evolution, in particular the developmental hourglass, is demonstrated. Finally, the possibility of collapse of dimensional reduction under nutrient limitation is discussed.
1 Introduction
Biological systems generally consist of a huge number of components, and are hierarchical: cells consist of diverse molecules, organisms consist of diverse cells, and ecosystems consist of diverse species[1]. In order to have robustness in a biological system, there must be some consistency across levels; for example, in order for a cell to grow, almost all the components in it must be produced in proportion to their abundances; otherwise the cellular state is not robust. In the field of universal biology that we intend to establish, we seek to uncover universal laws in adaptation, development, and evolution by noting robustness as well as plasticity to adapt to environmental changes. Here, we briefly review recent advances in the field, focusing on the evolutionary dimensional reduction of high-dimensional biological states, as well as the correlation between the changes in evolution and adaptation or development. Last, possible violation of the dimension reduction is discussed in relationship with the transition to dormant (non-growing) state of microorganisms.
2 Dimension reduction as a consequence of robustness
Recalling the diversity of components in a biological system, its state is high-dimensional: in a cell there are typically thousands of components. Then, the cellular state to be described by the concentration of each component requires a quite high-dimensional space; for a bacterial cell, about a few thousand protein or mRNA species (besides more metabolites), and the dimension of the state space exceeds at least a few thousand. Recent experimental and simulation studies, however, suggest that a drastic reduction in the dimensionality of the state change (phenotypes) in response to environmental changes has taken place[2-7].
First, it is noted that the steady state imposes some constraint on the phenotypic changes[2], but it alone cannot explain the experimentally observed dimensional reduction in the changes of many components across different environmental changes.
Then, we simulated a cell model consisting of a large number of components with catalytic-reaction network [3], in which phenotypes are generated as a result of intracellular reaction dynamics, where the abundances of each component change with catalytic reactions by taking nutrient chemicals from environment, leading to the cellular growth. When we adopted arbitrarily chosen reaction dynamics in the model, such reduction was not observed, even when the steady-growth condition of the cells was satisfied. In contrast, when these model cells were evolved to achieve a higher cell growth under a given environmental condition, it is found that the phenotypic changes of the cells upon a variety of environmental perturbations are constrained within a low-dimensional space.
The origin of such dimensional reduction has been explained as follows[3, 7]: Consider high-dimensional dynamical systems that shape an adapted phenotype. In general, such dynamics are subject to various perturbations. In fact, intracellular reactions are noisy because the number of molecules inside is not so large, and the environmental conditions also fluctuate. Then the fitted phenotype state should be robust to such perturbations.
Accordingly, the fitted state (attractor) should be attracted from most directions in the high-dimensional state space. In terms of dynamical systems, this implies that there is a strong attraction from most directions to the stationary state corresponding to the fitted state. However, if this strong attraction occurred from all directions, evolution to increase the fitness would be difficult because the state would not be easily changed by genetic mutations, which slightly alter the dynamical system (the flow in the state space). Then, along the direction of evolution, the phenotypes should be feasible to be changed. Accordingly, dominant changes in phenotypes (e.g., concentration of proteins) will be constrained along a one-dimensional curve (or few-dimensional space) in which evolution has progressed and will pregress (see Fig. 1).
In fact, we confirmed the above behavior in the evolution of a cell model with catalytic reaction dynamics by considering the relaxation dynamics to the steady state. By linearizing the dynamics around the steady state, one (or few) eigenvalues for the relaxation matrix (Jacobi matrix) is close to zero, as distinct from other eigenvalues which are much more negative [6].
Besides the cell model with catalytic reaction network, the evolutionary dimensional reduction has also been confirmed in the evolution of gene-regulatory network models with mutual activation or suppression of gene expressions[8], where fitness is determined by the expression levels of given target genes, depending on environmental conditions. In both models, the dynamics involve many variables that provide the high-dimensional phenotypic dynamics, while the fitness for evolution (e.g., the growth rate of a cell) is determined from these phenotypes. In these numerical studies, it was found that after evolution, changes in most variables induced by a variety of different environmental perturbations are highly constrained along a one-dimensional manifold.
Here, with the dimensional reduction, the relaxation dynamics towards the steady state consists of one (or few) slow modes. The evolution of such a slow mode is reasonable by the following argument: Consider a complex system with many interacting elements with similar time scales. It will be more difficult to evolve such system, because the random mutations to each element (gene) would mutually interfere. Then, directed phenotypic changes by genetic changes would be harder, as the proverb says too many cooks spoil the broth. In contrast, if few slow modes are separated, they will work for controlling others. In other words, the separation of a few slow modes, which is a result of evolution, also facilitates phenotypic evolution.
3 Relationship between environmental and evolutionary responses
Since the one-dimensional curve for the phenotypic changes in response to environmental changes follows the curve along which the evolution has progressed, the responses due to environmental and evolutionary (genetic) changes will occur along the same low-dimensional constraint (see Fig. 1b). Thus, phenotypic changes due to evolution are highly correlated with those due to environmental or stochastic perturbations, as confirmed by the simulation of the cell models.
Here let us denote the change in phenotype (e.g., the log concentration of each component) due to environmental responses as , and that due to genetic evolution as . Since all phenotype changes are then constrained along the common one-dimensional curve, and will follow, where is the change in the principal mode and is the angle between and . Then, will be independent over all components . ( is the case in which the principal axis is linear. If it is curved, one can use a function that represents the projection of with the principal axis, and the same result is obtained).
Now consider genetic evolution after applying an environmental stress that decreases the growth rate with . Evolution restores the growth rate so that the change from the original unstressed state, , should be less than . Here note that the growth rate of the cell is also a phenotype, and can be regarded as one of βs. Hence, the relationship
is expected to hold.
Since , this equation implies that the change in phenotype (e.g., logarithm of the concentrations of a protein or an mRNA species) induced by the stress is reduced as evolution progresses under this environmental stress, as . Thus, the long-term evolutionary response is to counteract the environmentally induced change. This homeostasis by adaptation is akin to the Le Chatelier principle in thermodynamics[4, 7]. In fact, eq. (1), together with the Le Chatelier-type responses, have been confirmed both in the simulation of cell models and in the evolution experiment of bacteria[4, 5].
4 Proportionality between the phenotypic variances by noise and by mutation
As mentioned above, intracellular dynamics (or any biological process that shapes the phenotype) are rather noisy, and phenotypic fluctuations are often rather large in cells[12, 13, 14]. Here, it should be recalled that in statistical physics, response and fluctuations are two sides of the same coin, as demonstrated by the fluctuation-response relationship[10, 11]. In fact, we have previously shown an evolutionary fluctuation-response relationship in which the evolutionary response (speed) is proportional to the variance of the phenotype over isogenic individuals (clones) induced by noise, denoted as (isogenic phenotypic variances)[15].
In considering the phenotypic variance, however, there is another type of fluctuations which are due to genetic changes (see FIg.2). Indeed, in standard population genetics, the genetic variance , which is the variance of the phenotype due to the genetic distribution, is known to be proportional to the rate of evolution[17, 18]. Accordingly, the proportionality between the phenotypic variances due to noise and due to genetic change is proposed, which is confirmed in the numerical simulation of the cell model over the course of evolution[19, 20].
Now, recall the proportionality of phenotypic responses by external environment and by genetic variation across components. We can then expect the proportionality of two phenotypic variances of by noise and by mutation, across components . This can be shown as follows: We denote the variance of the former by and that of the latter by . As shown in Fig. 1b, changes due to noise and genetic variation progress along the common principal axis. Thus, phenotypic variances due to the former are given by and those due to the latter by , where and are the average over the distribution of phenotypes induced by noise and mutations, respectively. Recall that the relaxation of phenotypic changes is much slower along the principal axis in Fig.1, phenotypic fluctuations due to noise are nearly confined to this axis. Thus, the variance of each component is given by the mean of the variance in the variable (along the principal axis) as (due to noise) and (due to mutation). Hence . and we obtain
will hold over the components.
This proportionality over most of the components was confirmed by using the aforementioned cell model with catalytic reaction network [4], as well as in the evolution of the gene regulatory network model[16, 21], after the evolution has proceeded to achieve robustness,
Note that this result introduces the possibility of predicting which component is likely to evolve: changes due to noise itself are not inherited, but genetic variation is naturally transmitted to offspring, and the resulting phenotypic variance is proportional to the evolutionary rate of that component [17, 18]. However, since is proportional to , the rate of evolution of is proportional to . In other words, as the variance of the component (trait) due to noise or environmental change is larger in prior to the genetic change, the component (trait) is more feasible to evolve. In other words, the direction in which it is likely to evolve in phenotypic space can be predicted in prior to the genetic evolution[7].
5 Summary of evolutionary dimensional reduction and - law
In summary, we have demonstrated the evolutionary dimensional reduction and 2-by-2 global proportionality of phenotypic changes between responses and fluctuations, and between perturbations due to environmental (noise) and genetic changes, as schematically shown in Fig. 3.
We expect that this evolutionary dimension reduction and the resulting constraint on phenotypic evolution will be universal in evolved biological systems. In the theoretical argument and cell models, the prerequisites for the emergence of dimension reduction are only the following two: (i) phenotypes with higher fitness are shaped by complex high-dimensional dynamical systems, which involve stochasticity and (ii) the fraction of such fit states is rare in the state space and in the genetic-rule space. These two features are necessary for our theoretical argument, while most biological systems satisfy them. Intuitively, a state that satisfies macro-micro consistency must be rare, and once it is achieved, its conservation is favored, implying the requirement of robustness.
6 Universality
Universality of evolutionary dimensional reduction, correlation between environmental and genetic changes, and the proportionality between genetic variance () and isogenic phenotypic variation () has recently been examined in theoretical models and experiments. Here we briefly review these.
(I) Reaction-network and gene-regulation network models:
The dimensional reduction by evolution has also been observed in several models. First, in reaction network models, phenotypic change is mainly constrained along a one-dimensional manifold across different types of fitness for selection (e.g., by selecting the concentration of a component). Second, even when the environmental conditions (e.g., the concentration of external nutrients) are not fixed but fluctuate over generations, the constraint within a one-dimensional manifold is observed [6]. Third, by using the gene expression dynamics model and by imposing a fitness condition depending on the expression pattern of target genes, the dimensional reduction and the relationship between variances are confirmed[8].
(II) Evolution of protein structure:
Through a combination of data analysis, elastic network models, and theoretical analysis, we have confirmed the relationship between and , as well as the dimensional reduction of protein structure. Changes in estimated structure due to conformation dynamics under noise and genetic mutations are constrained to a common low-dimensional space, typically about five dimensions, implying a strong reduction from the hundreds of residues that give rise to the high-dimensional state space [22]. A theoretical explanation is proposed there, while the consequence of few slow modes has been discussed earlier [23], and some model studies also suggest such dimensional reduction [24].
In relationship, we have also studied the evolving spin-Hamiltonian model. Here the spin variable takes up (+1) or down (-1) values, whose dynamics is governed by the interaction with other spins, with the energy , for spin and . (If is positive (negative), the two spins tend to be aligned (opposite)). By formulating the above interaction using a Hamiltonian () for energy, with the distribution under temperature (), where fitness is determined by a particular configuration of target spins and the matrix undergoes evolution by mutation and selection [25]. Under an intermediate temperature range, the evolutionary dimensional reduction and the relationship are then numerically confirmed [26, 27].
(III) Experimental confirmations: Dimensional reduction was first observed in the transcriptome analysis of bacteria E coli, under different types of stress[2]. Support for equation (1) was obtained by the bacterial evolution under ethanol stress[4, 5]. In addition, transcriptome analysis oacross thousands of mRNA expression data showed that changes in bacterial responses under different antibiotics could be predicted by about seven principal components [9]. The cross-fitness after evolution under different types of antibiotics is shown to be correlated with the response in prior to evolution, as predicted by the theory[8]. Further support for such dimensional reduction comes from Raman analysis of cellular components [28].
The proportionality between genetic and isogenic phenotypic variances across genes is also consistent with measurement of yeast gene expressions. Landry et al.[29] made global measurements of isogenic expression variation and mutation-induced variance. Although they used βmutational varianceβ (i.e., the rate of diffusive spread of each gene expression level upon the introduction of mutations) instead of , a high correlation between the isogenic phenotypic fluctuation () and the mutational variance was observed across all genes in yeast (see also [30]). More recently, the quantitative analysis of the gene expression pattern in Japanese medaka (fish) [31] has suggested proportionality.
The application of the proportionality between variances theory is not limited to gene expressions, but can be applied to any other phenotypic trait. The correlation between the isogenic variance and the evolution speed was confirmed in the evolution experiment of RNA structure[32]. In a series of experiments, Stearns et al. [33] measured the isogenic variance of five life-history traits in Drosophila melanogaster, while also measured the genetic variance in the same five traits between different genetic lines. Interestingly, they observed proportionality between the isogenic variance of each trait and its genetic variance. Further support for the - proportionality can be found in the in the wing morphology of Drosophila (fly) [34, 35].
(IV) Related topics:
The dimensional reduction has also been observed in the development of C. elegans[36] and in ecological evolution[37]. One possible link to the dimensional reduction will be the sloppy parameter hypothesis of Sethna [38], which suggests that many parameters used in biological models are irrelevant.
Besides the issue in biophysics, dimension reduction is attracting much attention in learning in the brain [39] and neural networks [40], as well as in machine learning. In deep learning, the possible network configurations that can support a desired output are quite redundant, and reduction to the appropriate output space is achieved. This is similar to our case. The genetic space (e.g., the network configuration) is high-dimensional, so there can be many ways to achieve the appropriate phenotype. Even though the final selected states are constrained in low-dimensional space, the use of high-dimensional space is essential for both machine learning and evolution.
7 Homeorhesis in Development and Evolution-Development Relationship
A promising extension of dimensional reduction and relationship will be in the area of evolutionary development.
To consider the robustness of the developmental process, Waddington, in the 1950s, schematically represented the developmental process by the motion of a ball along a landscape of branching valleys, as shown in Fig. 4[41]. The ball descends into lower valleys corresponding to stable cell types, while the successive branching of the valleys represents the differentiation process. In the figure, the horizontal axis represents the cell state that gives rise to the cell type, and the depth axis represents a slow developmental time. This picture has been revived in recent years and has received much attention.
Here, the horizontal axis represents the cellular state, possibly represented by the gene expression pattern. However, since there are thousands of genes (or components) in a cell, the cellular state involves many degrees of freedom, as mentioned above, while the horizontal axis represents only one variable. So how can cellular states potentially be represented by only one (or a few) variables? This first question is again a question of dimensional reduction, possibly achieved by evolution.
Since the stable state of each differentiated cell type is at the bottom of the valley, this may fit the attractor picture of a cellular state: For a given value of the depth axis , the cellular state is attracted to the bottom of the valley, where the state falls onto a fixed-point attractor, and no further change occurs. Here, however, the landscape itself is shaped along the depth axis with the developmental course, with which the valleys are successively shaped and deepened in a process known as canalization. Then the second question we need to address is: what does the y-axis represent for the (slower) landscape change and what is the origin of such a slow mode[42]?
Third, each developmental process is robust, i.e., how and when the valleys branch is rater robust. Waddington coined the term homeorhesis to discuss this robustness in the path[41]. This is not the stability of a state; how a path is robust for evolution is the third question that remains to be answered.
The fourth question concerns the correlation between evolution and development, which has been discussed since the time of Haeckel[43], who proposed the recapitulation theory, in which developmental processes follow a similar path to phylogenetic change. His original idea is not validated by itself, but the existence of some relationship between evolutionary and developmental changes has been noted for centuries [44, 45].
In this context, the concept of the developmental hourglass has recently gained much attention, in which different species within the same phylum increase their similarity in the mid-developmental stage. The difference between different species is large in the egg stage, then shrinks in the mid-developmental stage, and then increases again to have diverse adult forms. Because the bottleneck narrows in the mid-stage, it is called the hourglass hypothesis. When it was first proposed[46], it was difficult to determine quantitatively, but recently it has become possible to quantitatively compare gene expression profiles at each stage. The data seem to support the hypothesis [47, 48, 49], while further data analysis is needed to draw clear conclusions. Now, can we understand the developmental hourglass theoretically?
These four questions are now addressed in the light of the studies presented in Β§2-5, by extending the states to paths.
1) Dimensional reduction of developmental trajectories instead of evolutionary dimensional reduction of the state upon environmental changes.
2) Evolution of slow modes to control the developmental dynamics instead of evolution of slow modes for the intracellular process (gene expression dynamics).
3) Evolution of robustness in the developmental path rather than robustness of the final state.
4) Correlation between developmental and evolutionary changes, rather than correlation between the time course of phenotypic variances due to noise or environmental changes and phenotypic variation due to genetic changes, over developmental time .
To understand this evolutionary developmental relationship, we performed evolutionary simulations of the developmental process [50, 51] Specifically, cells follow the gene expression dynamics governed by the identical gene regulatory network. These cells are placed in one dimension, while some components are assumed to be exchanged with neighboring cells through diffusion interactions. Furthermore, external chemical components (morphogens) are supplied with a spatial gradient, providing input to the gene expression system. Under these conditions, this gene expression response and diffusion provides the dynamics for the gene expression level (protein concentration) in each cell. In the model, the gene expression levels of each cell reach a steady state that depends on the location of the cell, leading to pattern formation[50].
Let us then assume that the fitness is given by the number of final cell types defined by the expression pattern of target genes (specifically four target genes are adopted, so that cell types are possible according to the possibilities of all their on-off patterns.) By using the fitness and mutating the gene regulation networks over generations, the individuals with higher fitness are selected for the next generation[51].
After a few hundred generations, the networks have evolved to reach the fittest. It turns out that during evolution, the expression of a certain gene (protein) slowly changes, controlling the expression of the output genes. (For details, see the original paper[51]).
Now, the four questions raised above are resolved by the four points mentioned above(see Fig.5).
A) By calculating the similarity between the spatial patterns of gene expression levels of two species that evolved from a common ancestor, a developmental hourglass was observed, i.e., the difference in gene expression patterns between the two species is minimal in the middle stage of development. This mid-stage bottleneck (or peak of similarity between the two) is prominent when the species being compared are phylogenetically close, while it becomes vague and eventually disappears as the species diverge further in the phylogeny.
B) In addition to this developmental hourglass observation, other experimentally reported features for developmental hourglass were confirmed in the model. In particular, the variances of gene expression patterns across clones are minimal at the developmental bottleneck stage: The phenotypic variances due to initial or developmental noise decrease from the egg stage to the hourglass bottleneck stage, followed by an increase that leads to diversification in later developmental processes.
C) Behind the emergence of the developmental hourglass in the model, there exists the acquisition of genes which govern a slow scheduling process in development and induce developmental robustness. Such a slow mode controls the expression dynamics of other genes, which also facilitates evolution. In fact, these slow expression dynamics were conserved among organisms within a certain phylogenetic range of species sharing the hourglass (point 2).
D) Thus, the developmental hourglass can be related to the robustness and dimensional reduction by a slow mode as discussed in Β§2: The phenotypic diversity inherited from the parents at the egg stage must be reduced at some point, in order for the developmental process to proceed stably (point 3). This leads to the low-dimensional constraint on the developmental path of phenotype (point 1), leading to the decrease in variance at mid-developmental stage, the timing of which is given by this slowly changing gene expression. By presuming that it is correlated with the the the variance over mutants through the developmental time. the developmental bottleneck over species will be expected (point 4)[52].
8 Collapse of dimensional reduction: Anna Karenina phenomena?
So far we have discussed the achievement of dimensional reduction in a biologically good state. Then, it will be of interest to study if the reduction remains to be valid for cells with a bad, non-growing state which are put into poor environmental conditions. In fact, in bacteria, there generally exists a transition to the dormant state, i,e,, the growth-arrested state, when the nutrient concentration is decreased[53, 54], whereas theoretical models for the transition from an exponential growth state to a quiescent state (which cannot grow) have also been proposed[56, 55]. Will the dimensional reduction collapse with this transition to the dormant state?
In relation, a simple cell model that shows such transition to dormancy has recently been studied. The model consists of a catalytic reaction network with many components, where two-step elementary reaction processes with the formation of an intermediate complex are introduced[57], such as
If the rate of the second reaction is much faster, the product is formed as soon as the complex is formed, so that the above reaction processes are reduced to single mass action kinetics as , without intermediate complex formation. Then the model is reduced to the cell model mentioned already in Β§2-6. On the other hand, if the rate is not so fast, the intermediate complex remains formed for some time, and the complexes may accumulate depending on the condition. If such accumulation occurs, the abundances of free reactants that are not bound in complexes will decrease, which can hinder the reaction processes.
By numerically solving the above model with the production rate not so large, it is found that the growth rate decreases by orders of magnitude when the supplied nutrient concentration is reduced below a critical value, thus demonstrating the transition from exponential growth to the dormant phase[57].
This suppression of growth at the transition can be understood as follows: In the complex reaction network with many components, there is a set of components that catalyze each other for cell growth, i.e., an autocatalytic subnetwork, while some other components attached to it are produced with the help of the autocatalytic subnetwork. In the exponential growth phase, the abundances of the components are concentrated on the autocatalytic subnetwork, leading to the dimensional reduction. However, when the nutrient supply is reduced, the reaction flux is reduced and the complexes begin to accumulate. The catalytic reaction then stagnates in the pathways within the autocatalytic network, increasing the fraction of flux flowing to other components.
In fact, an increase in the diversity of chemical components is observed as the cell enters a non-growing quiescent state. This increase in diversity suggests a violation of dimensional reduction in an exponential growth state. In a bad state, the dimensional reduction achieved for a robust state with consistency between macro and micro levels may collapse, leading to diversity in components. Probably, the first sentence of Tolstoyβs Anna Karenina, βHappy families are all alike; every unhappy family is unhappy in its own wayβ, can be applied to cells or biological systems.
9 Summary
In this paper, we have reviewed recent advances in evolutionary dimension reduction. Noting that the biological system is robust to perturbations while maintaining plasticity for evolution, the phenotypic changes, which can originally be high-dimensional, are constrained to low-dimensional space. The responses to environmental changes, stochastic perturbations, and genetic changes follow dominantly along this low-dimensional manifold, and thus the responses to environmental and genetic changes, fluctuations to noise, and to genetic mutations are strongly correlated, allowing the prediction of phenotypic evolution. The extension of the dimensional reduction to the development-evolution relationship is briefly reviewed. Finally, a possible collapse of the reduction in a dormant (non-growing) cellular state is briefly discussed.
This research was partially supported by and the Novo Nordisk Foundation (NNF21OC0065542) and Grant-in-Aid for Scientific Research (A) (20H00123) from the Ministry of Education, Culture, Sports, Science and Technology (MEXT) of Japan. The author would like to thank Chikara Furusawa, Takahiro Kohsokabe, Takuya Sato, Ayaka Sakata, Qian-Yuan Tang, and Tetsuhiro S. Hatakeyama for useful discussions.